# Learning to Animate Images from A Few Videos to Portray Delicate Human Actions

### v1

@@ Abstract
Despite recent progress, video generative models still struggle to animate human actions from static images, particularly when handling uncommon actions whose training data are limited. In this paper, we investigate the task of learning to animate human actions from a small number of videos—16 or fewer—which is highly valuable in real-world applications like video and movie production. Few-shot learning of generalizable motion patterns while ensuring smooth transitions from the initial reference image is exceedingly challenging. We propose FLASH (Few-shot Learning to Animate and Steer Humans), which improves motion generalization by aligning motion features and inter-frame correspondence relations between videos that share the same motion but have different appearances. This approach minimizes overfitting to visual appearances in the limited training data and enhances the generalization of learned motion patterns. Additionally, FLASH extends the decoder with additional layers to compensate lost details in the latent space, fostering smooth transitions from the initial reference image. Experiments demonstrate that FLASH effectively animates images with unseen human or scene appearances into specified actions while maintaining smooth transitions from the reference image. The animated videos can be accessed on the webpage https://lihaoxin05.github.io/human_action_animation/.
@@ 1 Introduction
Despite substantial progress [27, 56, 82, 21, 65, 9, 76, 40, 78, 22, 69, 66, 64], video generative models still struggle to accurately portray human actions from static images. Even commercial AI video generators, such as Dream Machine [^1] https://lumalabs.ai/dream-machine from Luma AI and KLING AI [^2] https://www.klingai.com/ from Kuaishou, encounter difficulty with this task. As shown in Figure 1, both models fail to animate actions such as balance beam jump or shooting a soccer ball from static images. This difficulty arises from the scarcity of training data that specifically depict the target action. As human actions are diverse and likely follow a long-tailed distribution, many highly recognizable human actions, such as those of a niche sport like balance beam, suffer from limited training data. The data scarcity prevents data-hungry video generative models from effectively learning such actions.
@@
Figure 1: Comparison of animated human action videos produced by Dream Machine, KLING AI, and FLASH (our method). In the balance beam jump action, Dream Machine produces unrealistic, physics-defying movements, whereas KLING AI generates a jump but fails to portray standard jumps on the balance beam. For the soccer shooting action, both Dream Machine and KLING AI struggle to generate the correct shooting motion and the person never kicks the ball away. In contrast, FLASH successfully animate actions that resemble the real-world actions in the last row. We provide additional examples in Figure A1 of Appendix.
In this paper, we explore the task of learning to animate human actions from a small set of videos. Our aim is to transform a static reference image into a short video of a few seconds, which portrays a specific human action described by a textual prompt. This transformation is learned from a limited dataset containing up to 16 videos for each action class, thereby reducing the need for extensive video data collection. This capability holds the promise to reduce computational cost and broaden the application domains of video generative models; it is particularly valuable for applications like video and movie production, which needs to animate specific actors performing a wide range of actions, yet each action is only used once or twice. Under such use cases, techniques requiring many example videos for each action become cost-ineffective.
@@
Existing image animation methods encounter considerable difficulties with this task. These approaches typically rely on large video datasets for training and primarily focus on preserving the appearance of the reference images [72, 19, 32, 15, 63, 20, 44, 52, 15, 80] or on learning spatial-temporal conditioning controls (e.g., optical flows) to guide image animation [47, 33, 55]. However, these methods become impractical for the few-shot task. When limited to no more than 16 videos, these methods suffer from severe overfitting and fail to learn generalizable motion patterns and object transformations. [68, 81] employ a two-path approach to customize motion from a few videos, but they require training for each reference image for animation, leading to limited flexibility. Although [46, 70, 34, 37] attempt to learn appearance-irrelevant motion patterns from limited data, their models lack explicit supervision for appearance-general motion, which limits performance.
The main challenge of this few-shot task is learning generalizable motion patterns. The limited number of training videos makes it difficult to learn motion patterns that generalize to diverse appearances. In addition, the initial reference image adds an extra condition, requiring the motion to align with the spatial arrangement of humans or objects in the image to maintain smooth transitions from it. The few-shot learning of motion that generalize to user-provided reference images is more challenging.
@@
To tackle this challenge, we propose FLASH (Few-shot Learning to Animate and Steer Humans), a method for few-shot human action animation. To learn generalizable motion patterns, FLASH devise the Motion Alignment Module to align the motion features and inter-frame correspondence relations between a video and its strongly augmented variant, where the motion remains the same but the appearance differs significantly. By requiring the model to predict the two videos using the two aligned motion patterns, this approach encourages learning motion patterns that can generalize across different appearances, reducing overfitting to the appearance in the limited training data. Additionally, to improve transition smoothness from the reference image, FLASH employs the Detail Enhancement Decoder to propagate the details in the reference image to generated frames, which compensates for the loss of details in the latent space in the decoding process. The overall framework of FLASH is illustrated in Figure 2 (a).
Through experiments on 12 atomic human actions selected from HAA500 [7], we demonstrate that FLASH accurately animates human actions from diverse reference images while maintaining smooth transitions. It outperforms existing image animation methods across various quantitative metrics and human evaluations, showcasing the effectiveness and superiority of FLASH. It also generalizes to non-realistic figures like cartoon characters or humanoid aliens. Our contributions include: (1) We tackle the practical and challenging task of few-shot human action animation, an under-explored area with significant potential for video and film production. (2) We propose FLASH, a framework designed to learn generalizable motion patterns from limited training data. (3) Experiments on 12 atomic human actions validate the effectiveness of FLASH.
@@ 2 Related Work
Video Generation. Video generation using diffusion models [25, 58, 57] have notably surpassed methods based on GANs [16], VAEs [36] and flow techniques [5]. Diffusion models for video generation can be broadly classified into two groups. The first group generates videos purely from textual descriptions. These methods extend advanced text-to-image generative models by integrating 3D convolutions, 3D UNets, and temporal attention modules to capture temporal dynamics in videos [27, 26, 56, 82, 2, 21, 65]. To mitigate concept forgetting when training on low-quality videos, some methods use both videos and images jointly for training [27, 4]. Large Language Models (LLMs) contribute by generating frame descriptions [18, 28, 38] and scene graphs [11] to guide the video generation. Trained on large-scale video-text datasets [1, 73, 6], these methods excel at producing high-fidelity videos. However, they typically lack control over specific frame layouts, such as object positions and human poses. To improve controllability, LLMs are used to predict control signals [42, 39, 43], but these signals typically offer coarse control (e.g., bounding boxes) rather than fine-grained control (e.g., detailed human motion or object deformation).
@@
On top of textual descriptions, the second group of techniques benefit from additional guidance sequences, such as depth maps, motion vectors, optical flows, and bounding boxes [9, 76, 40, 78, 22, 66, 64], which help control motion and frame layouts. Additionally, several techniques use existing videos as guidance to generate videos with different appearances but identical motion patterns [69, 50, 74, 13, 75, 79, 41, 53, 49, 31]. However, these methods cannot create novel videos that share the same motion class with the guidance video but differ in the actual motion, such as human positions and viewing angles, which limits their generative flexibility.
@@
Image Animation. Image animation involves generating videos that begin with a given reference image. Common approaches achieve this by integrating the image features into videos through cross-attention layers [63, 72, 19, 32, 15], employing additional image encoders [20, 67], or incorporating the reference image into noised videos [77, 70, 14, 44, 52, 15]. Another line of methods focuses on learning structural guidance (e.g., motion maps) that aligns with the reference image to guide the generation of subsequent frames [55, 47, 33]. However, these approaches often require extensive training videos to learn motion or structure guidance. [81, 68] employ a temporal path to learn motion patterns from a few videos and a spatial path to learn appearance from a reference image for animation. However, the need for training on each reference image restricts their flexibility. While [46, 70, 34, 37] are similar to our work in learning specific motion patterns from a few videos, they use the reference image as an appearance condition and rely on the model to automatically prioritize motion over appearance. Without explicit supervision for appearance-general motion, their generalizability is still limited. In this paper, we propose FLASH, which learns generalizable motion from only a few videos through explicit supervision for appearance-general motion, and the learned motion can be applied to reference images that differ widely in visual attributes like human positions and texture.
Figure 2: (a) Overview of the FLASH framework. To learn generalizable motion patterns, (b) the Motion Alignment Module aligns motion features and inter-frame correspondence relations (Corr. Rel.) between a training video and its strongly augmented version (see Sec. 3.2). To improve the transition smoothness from the reference image to subsequent frames, (c) the Detail Enhancement Decoder propagates hierarchical details from the reference image into the generated frames (see Sec. 3.3).
@@ 3 FLASH
To learn generalizable motion from a limited set of training videos while maintaining smooth transition from the reference image, we propose FLASH, which builds upon a video diffusion model (see Sec. 3.1) and features two novel components as illustrated in Figure 2. The first is the Motion Alignment Module, designed to learn robust motion patterns that generalize across different appearances, which will be detailed in Sec. 3.2. The second is the Detail Enhancement Decoder, which propagates details from the reference image to generated frames to enhance transition smoothness, and will be explained in Sec. 3.3.
@@ 3.1 Preliminaries
Image Diffusion Models. Latent Diffusion Models (LDM) [54], a leading image generative model, comprises four main components: an image encoder $\mathcal{E}$, an image decoder $\mathcal{D}$, a text encoder $\mathcal{T}$, and a U-Net $\epsilon_{\theta}$. During training, an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ is first encoded into a latent image $\bm{z}_{0}=\mathcal{E}(\bm{x})\in\mathbb{R}^{h\times w\times c}$, where $h$, $w$ and $c$ denote the height, width and number of channels of the latent image, respectively. Next, $\bm{z}_{0}$ undergoes a pre-defined diffusion process [8, 25] to add noise, resulting in $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\,I)$, $t\in[0,T]$ denotes the noising step, and $\bar{\alpha}_{t}$ denotes the noise strength. The U-Net is then trained to predict the noise $\bm{\epsilon}_{t}$ from $\bm{z}_{t}$. During inference, a latent noise $\bm{z}_{T}$ is drawn from $\mathcal{N}(\bm{0},\,I)$ and progressively denoised into $\hat{\bm{z}}_{0}$. Finally, the decoder reconstructs the generated image $\hat{\bm{x}}=\mathcal{D}(\hat{\bm{z}}_{0})$.
Video Diffusion Models. The LDM framework can be naturally extended to generate videos. Given a video consisting of $N$ frames $\bm{X}=\langle\bm{x}^{i}\rangle_{i=1}^{N}$, each frame is encoded into a latent frame $\bm{z}_{0}^{i}=\mathcal{E}(\bm{x}^{i})\in\mathbb{R}^{h\times w\times c}$. Collectively, all latent frames form a latent video $\bm{Z}_{0}=\langle\bm{z}_{0}^{i}\rangle_{i=1}^{N}\in\mathbb{R}^{N\times h\times w\times c}$ used in the noising and denoising processes. The training loss is defined as:
$\mathcal{L}_{D}=\mathbb{E}_{\bm{X},\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\,I),t,y}\left[\left\|\bm{\epsilon}_{t}-\epsilon_{\theta}\left(\bm{Z}_{t},t,\mathcal{T}(y)\right)\right\|^{2}_{2}\right],$ (1)
where $y$ is the text prompt associated with the video. To capture temporal dynamics in videos, temporal attention layers are integrated into the U-Net [27, 9, 21, 20]. To enhance consistency between frames, the self-attention layers in the U-Net are replaced with cross-frame attention layers [35, 70], in which features from the first frame (the reference frame) are used as the key and value, enabling the appearance of the first frame to be propagated to subsequent frames. In image animation tasks, the noise-free reference image is integrated into the noised latent video [70, 52] to preserve the appearance of the reference image. Further details are in Appendix Sec. A2.1. FLASH inherits all these components of a standard video diffusion model.
@@ 3.2 Motion Alignment Module
The Motion Alignment Module directs the model to learn motion that generalizes across various appearances. To achieve this, we force the model to learn consistent motion patterns from a pair of videos with identical motion but different appearances, created using strong data augmentation. We align two motion signals in the U-Net between the video pairs and requires the model to predict both videos using the same aligned motion signals. This encourages the model to learn general motion patterns across both videos rather than overfitting to appearance-specific details in each one. The overall process is depicted in Figure 2 (b) and elaborated in the following sections.
@@
Strongly Augmented Videos. From the original video, $\bm{X}^{\text{ori}}$, we create a strongly augmented version $\bm{X}^{\text{aug}}$, which has different appearances but the same motion information. Here we choose the augmentations as Gaussian blur with random kernel sizes and random color adjustments. Details of the augmentations and example augmented videos are in Appendix Sec. A2.2. The overall loss is the diffusion noise prediction, aimed to recover the two videos.
$\mathcal{L}_{D}=\mathbb{E}_{\bm{X}^{\text{ori}},\bm{X}^{\text{aug}},\bm{\epsilon}_{t},t,y}\sum_{\text{v}\in\{\text{ori},\text{aug}\}}\bigg{\|}\bm{\epsilon}_{t}^{\text{v}}-\epsilon_{\theta}\Big{(}\bm{Z}_{t}^{\text{v}},t,\mathcal{T}(y)\Big{)}\bigg{\|}^{2}_{2}.$ (2)
For simplicity, we omit the superscripts ori and aug when the same operation is applied to both videos.
Motion Feature Alignment. The purpose of motion feature alignment is to force the model to learn the same motion features from the videos before and after the strong augmentation, which distorts appearance but not motion. We require the model to recover the augmented video from motion features of the original video and the appearance features of the augmented video. This encourages learning of consistent motion features from both videos. We denote the features extracted after a temporal attention layer as $\bm{F}_{\text{in}}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$. Since motion is represented by the temporal changes of the features, we remove the static components from $\bm{F}_{\text{in}}$ and normalize it to obtain the dynamic features:
$\hat{\bm{F}}_{\text{in}}=\frac{\bm{F}_{\text{in}}-\bm{\mu}_{\text{T}}(\bm{F}_{\text{in}})}{\bm{\sigma}_{\text{T}}(\bm{F}_{\text{in}})},$ (3)
where $\bm{\mu}_{\text{T}}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c^{\prime}}$ and $\bm{\sigma}_{\text{T}}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c^{\prime}}$ are the mean and standard deviation of $\bm{F}_{\text{in}}$ calculated along the temporal dimension. The standard deviation serves as a normalization factor to reduce the influence of feature scales (e.g., varying brightness in videos). As a result, $\hat{\bm{F}}_{\text{in}}$ becomes independent of static appearance elements and is focused on the changes within the video.
However, motion information is predominantly encoded in a few channels [71], and we need to identify the channels with rich motion information. We quantify the motion information using the standard deviations along the temporal dimension in each channel, which are then averaged across all spatial positions, and the result is denoted as $\bm{s}\in\mathbb{R}^{c^{\prime}}$. Channels whose value in $\bm{s}$ exceed the $\tau$-percentile are identified as motion channels and denoted as the set $\mathcal{C}^{m}$. The motion features are thus represented as $\hat{\bm{F}}_{\text{in}}[c],\forall c\in\mathcal{C}_{m}$.
We denote the motion features of the original video as $\hat{\bm{F}}^{\text{ori}}_{\text{in}}[c]$, and those of the augmented video as $\hat{\bm{F}}^{\text{aug}}_{\text{in}}[c]$. We replace $\hat{\bm{F}}^{\text{aug}}_{\text{in}}[c]$ with $\hat{\bm{F}}^{\text{ori}}_{\text{in}}[c]$ as follows:
$\hat{\bm{F}}^{\text{aug}}_{\text{out}}[c]\leftarrow\hat{\bm{F}}^{\text{ori}}_{in}[c],\quad\forall c\in\mathcal{C}_{m}.$ (4)
Finally, we restore the features with video-specific mean and standard deviation, $\bm{F}_{\text{out}}^{\text{ori}}=\hat{\bm{F}}_{\text{out}}^{\text{ori}}\bm{\sigma}_{\text{T}}^{\text{ori}}+\bm{\mu}_{\text{T}}^{\text{ori}},\bm{F}_{\text{out}}^{\text{aug}}=\hat{\bm{F}}_{\text{out}}^{\text{aug}}\bm{\sigma}_{\text{T}}^{\text{aug}}+\bm{\mu}_{\text{T}}^{\text{aug}}$, which are used in noise prediction $\epsilon_{\theta}(\cdot)$.
Inter-frame Correspondence Relation Alignment. The purpose of inter-frame correspondence relation alignment is to guide the model to learn the same cross-frame motion between the original and augmented videos. From the attention weights of the original video, we identify spatial correspondence between the first frame and subsequent frames. We require the reconstruction of the augmented video to adopt the same spatial correspondence. This forces the model to learn a general warping strategy across both videos. Since the video pairs have the same motion but different appearance, the learned warping strategy becomes motion-sensitive and appearance-invariant.
@@
We denote the input features of a cross-frame attention (CFA) layer as $\bm{F}_{\text{in}}=\langle\bm{f}^{i}_{\text{in}}\rangle_{i=1}^{N}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$. The output features are computed as:
$\bm{F}_{\text{out}}=\text{Softmax}\left(\frac{(\bm{Q}\bm{W}^{Q})(\bm{K}\bm{W}^{K})^{\top}}{\sqrt{c^{{}^{\prime}}}}\right)(\bm{V}\bm{W}^{V})=\bm{S}(\bm{V}\bm{W}^{V}),$ (5)
@@
where $\bm{Q}=\bm{F}_{\text{in}}$, $\bm{K}=\bm{f}^{1}_{\text{in}}$, $\bm{V}=\bm{f}^{1}_{\text{in}}$ are the query, key, and value, respectively, and $\bm{W}^{Q}$, $\bm{W}^{K}$, $\bm{W}^{V}$ are the learnable projection matrices. The key and value are from the first frame of the video. Therefore, $\bm{S}$ represents the similarity between the query and the key from the first frame, which implicitly warps the first frame into subsequent frames [45]. Consequently, $\bm{S}$ can be interpreted as correspondence relations between spatial locations of the first frame and those of subsequent frames, capturing cross-frame motion.
We denote the inter-frame correspondence relations of the original video and the augmented video as $\bm{S}^{\text{ori}}$ and $\bm{S}^{\text{aug}}$. We replace $\bm{S}^{\text{aug}}$ with $\bm{S}^{\text{ori}}$ in the network processing the augmented video. Effectively, this amounts to using $\bm{S}^{\text{ori}}$ to warp the features of the first frame of the augmented video to produce outputs $\bm{F}_{\text{out}}^{\text{aug}}$, which are used to reconstruct the augmented video in $\epsilon_{\theta}(\cdot)$.

### v2

@@ Abstract
Despite recent progress, video generative models still struggle to animate static images into videos that portray delicate human actions, particularly when handling uncommon or novel actions whose training data are limited. In this paper, we explore the task of learning to animate images to portray delicate human actions using a small number of videos—16 or fewer—which is highly valuable for real-world applications like video and movie production. Learning generalizable motion patterns that smoothly transition from user-provided reference images in a few-shot setting is highly challenging. We propose FLASH (Few-shot Learning to Animate and Steer Humans), which learns generalizable motion patterns by forcing the model to reconstruct a video using the motion features and cross-frame correspondences of another video with the same motion but different appearance. This encourages transferable motion learning and mitigates overfitting to limited training data. Additionally, FLASH extends the decoder with additional layers to propagate details from the reference image to generated frames, improving transition smoothness. Human judges overwhelmingly favor FLASH, with 65.78% of 488 responses prefer FLASH over baselines. We strongly recommend watching the videos in the Webpage [^1] https://lihaoxin05.github.io/human_action_animation/, as motion artifacts are hard to notice from images.
@@ 1 Introduction
Despite substantial progress [33, 75, 110, 26, 86, 11, 99, 50, 102, 28, 91, 87, 85, 98, 45], video generative models still struggle to accurately portray delicate human actions, especially when they are required to start from a user-provided reference image. Even commercial AI video generators trained on large-scale datasets, such as KLING AI [^2] https://www.klingai.com/image-to-video and Wanx AI [^3] https://tongyi.aliyun.com/wanxiang/videoCreation, encounter difficulty with this task. As shown in Figure 1, both fail to animate actions such as balance beam jump or shooting a soccer ball.
We attribute the difficulty mainly to the interplay between action complexity and training data scarcity. Due to skeletal and joint structures, human actions are highly complex and constrained in unique ways. For example, the range of forearm pronation is typically 80-90°, whereas the neck can do 150° rotation and 125° flexion [5]. Inferring such precise specifications from videos alone is a difficult inverse problem. To make things worse, human action videos follow a long-tailed distribution [104], which means that, for a wide range of human actions, only a small amount of data are available for learning. The action recognition community has long recognized this problem and created a body of literature on few-shot and open-vocabulary action recognition [101, 65, 37, 51]. Interestingly, few-shot generation of human actions remains under-investigated.
@@
Figure 1: Comparison of animated human action videos produced by KLING AI, Wanx AI and FLASH (our method). In the balance beam jump action, Wanx AI produces physics-defying movements, whereas KLING AI generates a jump but fails to portray the standard jump on the balance beam. For the soccer shooting action, both KLING AI and Wanx AI struggle to generate the correct shooting motion and the person never kicks the ball away. In contrast, FLASH successfully generates actions that resemble the real-world actions in the last row. We strongly recommend watching the animated videos in the Webpage, as motion artifacts can be hard to notice from static images.
In this paper, we explore the task of animating an image to portray delicate human actions by learning from a small set of videos. The problem input includes a static reference image as well as a textual prompt describing the action. We learn from up to 16 videos for each action class, thereby reducing the need for extensive video data collection. The animation should begin exactly with the reference image, giving users precise control over the initial state of the video. This capability is particularly valuable for applications like video and movie production, which need to animate diverse actors performing uncommon or newly designed actions, often with only a few example videos available and requiring precise control over the initial states of the action, such as the actors’ orientations and spatial arrangement of the scenes. Controllable generation has been studied for both images [63, 74, 103] and videos [27, 99, 55], but the additional constraint also renders the problem more difficult.
@@
Existing image animation methods encounter considerable difficulties with this task. These approaches typically rely on large video datasets for training and primarily focus on preserving the appearance of the reference images [94, 23, 40, 18, 84, 25, 56, 68, 18, 106, 24] or learning spatial-temporal conditioning controls (e.g., optical flows) to guide image animation [60, 41, 73]. However, with limited training data, these methods suffer from overfitting and struggle to learn generalizable motion patterns. Customized video generation methods [58, 90, 108, 47] can learn target motion from a few examples but fall short in the ability to align the action with the user-provided reference image.
@@
We propose FLASH (Few-shot Learning to Animate and Steer Humans), a framework for few-shot human action animation. To learn generalizable motion patterns from limited videos, FLASH introduces the Motion Alignment Module, which forces the model to reconstruct a video using the motion features of another video with the same motion but different appearance. This facilitate the learning of generalizable motion features and reduces overfitting. Additionally, FLASH employs a Detail Enhancement Decoder to propagate multi-scale details from the reference image to the generated frames, providing smooth transition from the first reference frame.
Experiments on 16 delicate human actions demonstrate that FLASH accurately and plausibly animates diverse reference images into lifelike human action videos. Human judges overwhelmingly favor FLASH; 65.78% of 488 responses prefer FLASH over open-source baselines. FLASH also outperforms existing methods across automatic metrics, and generalizes to non-realistic figures like cartoon characters or humanoid aliens.
@@ 2 Related Work
Video Generation. Video generation using diffusion models [31, 78, 77] have notably surpassed methods based on GANs [19], VAEs [44] and flow techniques [6]. Diffusion models for video generation can be broadly classified into two groups. The first group generates videos purely from text descriptions. These methods extend advanced text-to-image generative models by integrating 3D convolutions, temporal attention layers, or 3D full attention layers to capture temporal dynamics in videos [33, 32, 75, 110, 2, 26, 86, 98, 45]. To mitigate concept forgetting when training on videos, some methods use both videos and images jointly for training [33, 4, 45]. Large Language Models (LLMs) contribute by generating frame descriptions [21, 35, 48] and scene graphs [13] to guide the video generation. Trained on large-scale video-text datasets [1, 95, 7], these methods excel at producing high-fidelity videos. However, they typically lack control over frame layouts like object positions. To improve controllability, LLMs are used to predict control signals [53, 49, 54], but these signals typically offer coarse control (e.g., bounding boxes) rather than fine-grained control (e.g., human motion or object deformation).
@@
On top of text descriptions, the second group of techniques uses additional control sequences, such as depth maps, optical flows, trajectories and bounding boxes [11, 99, 50, 102, 28, 107, 87, 85, 89, 34, 55], to control frame layouts and motion. Additionally, several techniques use existing videos as guidance to generate videos with different appearances but identical motion [91, 66, 96, 16, 97, 105, 52, 69, 64, 39, 93]. However, these methods cannot create novel videos that share the same motion class with the guidance video but differ in the actual motion, such as human positions and viewing angles, which limits their generative flexibility.
@@
Image Animation. Image animation involves generating videos that begin with a given reference image controlling the initial action states. Common approaches achieve this by integrating the image features into videos through cross-attention layers [84, 94, 40, 18], employing additional image encoders [25, 23, 88], or incorporating the reference image into noised videos [100, 92, 17, 56, 68, 18]. Another line of methods focuses on learning guidance sequences (e.g., motion maps) that aligns with the reference image to guide the generation of subsequent frames [73, 60, 41]. However, these approaches often require extensive training videos to learn motion or guidance sequences, making them ineffective with limited data.
Customized Generation. Customized generation creates visual content tailored to specific concepts using limited samples. In the image domain, static concepts are associated with new texts [14, 71, 72, 46] or model parameters [8, 76, 22]. In the video domain, [59, 58, 90, 108, 47] learn target appearance and motion from limited data but lack control over initial action states, making it difficult for users to control the positions and directions of the actor and objects. Additionally, their requirement for test-time training on each reference image limits flexibility. While [92, 42] are similar to our work in learning specific motion patterns from a few videos, they rely on the model to automatically prioritize motion over appearance, which limit the generalizability due to the lack of explicit guidance for appearance-general motion. In contrast, our work learns generalizable motion from a few videos with explicit guidance, enabling it to generalize to reference images with varying visual attributes, such as actor positions and textures.
Figure 2: An illustration of the Motion Alignment Module. Both the noised latent representations of the original and strongly augmented videos are input to the U-Net. In the temporal attention layers, static and motion features are extracted from both videos. Motion features from the original video are transferred to the augmented video (red arrows), and the recombined features are passed to the next layer. In the cross-frame attention layers, attention scores from the original video, which capture its cross-frame motion structure, are used to warp the augmented video (red arrow) before passing it to the next layer. The U-Net is trained to predict the noise added to both videos based on the motion patterns of the original video, encouraging the learning of consistent motion patterns.
@@ 3 FLASH
Building upon latent video diffusion (Sec. 3.1), we propose two novel system components. The first is the Motion Alignment Module, detailed in Sec. 3.2, which encourages the learning of generalizable motion patterns and prevents overfitting to static features. The second is the Detail Enhancement Decoder, explained in Sec. 3.3, which propagates details from the user-provided reference image to generated frames to enhance transition smoothness.
@@ 3.1 Preliminaries
Latent Image Diffusion Model. Latent Diffusion Model (LDM) [70] comprises four main components: an image encoder $\mathcal{E}$, an image decoder $\mathcal{D}$, a text encoder $\mathcal{T}$, and a U-Net $\epsilon_{\theta}$. The training begins with an image $\bm{x}$ and a textual description $y$. We encode the image into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})\in\mathbb{R}^{h\times w\times c}$. Next, for a randomly sampled time index $t$, we add Gaussian noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\,I)$ to the latent, yielding a noised version $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t}$, where $\bar{\alpha}_{t}$ is the noise strength [10, 31]. The main step is to train the U-Net $\epsilon_{\theta}(\bm{z}_{t},t,\mathcal{T}(y))$ to predict the added noise $\bm{\epsilon}_{t}$, so that it can be subtracted to recover the original latent $\bm{z}_{0}$. Since identifying the noise and identifying the original latent representation are two sides of the same coin, we often refer to the task of the U-Net as reconstructing the original latent representation. During inference, we randomly sample a latent noise $\bm{z}_{T}\sim\mathcal{N}(\bm{0},\,I)$ and progressively perform denoising to obtain an estimated noise-free latent image $\hat{\bm{z}}_{0}$ using the trained U-Net. Finally, the decoder recovers the generated image in pixel space $\hat{\bm{x}}=\mathcal{D}(\hat{\bm{z}}_{0})$.
Latent Video Diffusion Model. The LDM framework can be naturally extended to video generation. Given a video consisting of $N$ frames $\bm{X}=\langle\bm{x}^{i}\rangle_{i=1}^{N}$, we encode each frame and yield a latent video representation $\bm{Z}_{0}=\langle\bm{z}_{0}^{i}\rangle_{i=1}^{N}\in\mathbb{R}^{N\times h\times w\times c}$. The training is the same as before with the following loss:
$\mathcal{L}_{D}=\mathbb{E}_{\bm{X},y,\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\,I),t}\left[\left\|\bm{\epsilon}_{t}-\epsilon_{\theta}\left(\bm{Z}_{t},t,\mathcal{T}(y)\right)\right\|^{2}_{2}\right].$ (1)
The video denoising U-Net differs from the image counterpart in a few ways. To capture temporal dynamics, [33, 11, 26, 25] add temporal attention after each spatial attention layer. To enhance consistency with the reference frame, [43, 92] replace spatial self-attention in the U-Net with spatial cross-frame attention, where features from the reference frame (typically the first frame) serve as the keys and values. These layers propagate appearance features from the reference frame to other frames to improve consistency. Further, the reference frame is kept noise-free in the noised latent video to preserve its appearance [92, 68]. FLASH incorporates all these components. Further details are in Appendix A2.1.
@@ 3.2 Motion Alignment Module
The Motion Alignment Module forces the model to learn motion patterns that remain consistent despite appearance changes. This is achieved by four steps. First, given a training video, we create a strongly augmented video with the same motion but different appearances. Second, we identify feature channels that capture motion information and static information from each video. We recombine the motion features of the original video and static features of the augmented video as the new features of the augmented video. Third, we force the cross-frame spatial attention in the augmented video to adhere to the attention in the original video, which aligns the motion structures of the two videos, as reflected by the attention weights. Finally, we train the network to reconstruct the latent (i.e. predicting the added noise) of the augmented video using the the motion features and structures of the original video, which encourages the learned motion to be generalizable across different appearances. The overall process is depicted in Figure 2 and elaborated below.
@@
Strongly Augmented Videos. Given an original video, $\bm{X}^{\text{ori}}$, we create a strongly augmented version $\bm{X}^{\text{aug}}$, which has different appearances but the same motion. We choose the augmentations as Gaussian blur with random kernel sizes and random color adjustments. The randomness of augmentation ensures that the model encounters different original-augmented video pairs at different training epochs. Details of the augmentations and example augmented videos are in Appendix A2.2.
Identifying Motion Channels. We set out to identify motion features in the U-Net from the original video. We denote the features extracted by a temporal attention layer as $\bm{F}_{\text{in}}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$, and compute its mean $\bm{\mu}_{\text{T}}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c^{\prime}}$ and standard deviation $\bm{\sigma}_{\text{T}}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c^{\prime}}$ along the temporal dimension. [93] shows that motion information is predominantly encoded in a few channels. Thus, we take the average of $\bm{\sigma}_{\text{T}}$ across spatial positions, denoted as $\bm{s}\in\mathbb{R}^{c^{\prime}}$, and consider channels in the top $\tau$-percentile as motion channels. We identify the channel indices from the original video and consider the corresponding channels in the augmented video also as motion channels.
Transferring Motion Channels. The next step is to transfer motion channels from the original video to the augmented video, in order to encourage the network learn the motion channels generalizable to both videos. Before transferring, we remove the static components from $\bm{F}_{\text{in}}$ by normalization:
$\hat{\bm{F}}_{\text{in}}=\frac{\bm{F}_{\text{in}}-\bm{\mu}_{\text{T}}(\bm{F}_{\text{in}})}{\bm{\sigma}_{\text{T}}(\bm{F}_{\text{in}})}.$ (2)
We posit that normalization by the standard deviation reduces the influence of feature scales (e.g., varying brightness) but preserves motion.
After that, with the motion channels identified from the original video, we replace the corresponding channels in the augmented video with those from the original, yielding a new feature map $\hat{\bm{F}}^{\text{aug}}_{\text{out}}$. Finally, we restore the video mean and standard deviation of the augmented video, $\bm{F}_{\text{out}}^{\text{aug}}=\hat{\bm{F}}_{\text{out}}^{\text{aug}}\bm{\sigma}_{\text{T}}^{\text{aug}}+\bm{\mu}_{\text{T}}^{\text{aug}}$, which are fed to the next layer and are eventually used in noise prediction. The original video features remain unchanged and are unaffected by this step.
Cross-frame Attention Alignment. The purpose of this technique is to guide the model to learn the same cross-frame motion structures from the two videos. Recall that the cross-frame attention treats the reference frame (the first frame) as the keys and values, and the current frame as the queries. Hence, the attention weights indicate how patches in the reference frame correspond to patches in the current frame. We want these correspondences to be identical in both videos.
@@
We denote the input features of a cross-frame attention layer as $\bm{F}_{\text{in}}=\langle\bm{f}^{i}_{\text{in}}\rangle_{i=1}^{N}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$. The output features are computed as:
$\bm{F}_{\text{out}}$ $=\text{Softmax}\left(\frac{(\bm{Q}\bm{W}^{Q})(\bm{K}\bm{W}^{K})^{\top}}{\sqrt{c^{{}^{\prime}}}}\right)(\bm{V}\bm{W}^{V})$ (3)
$=\bm{S}(\bm{V}\bm{W}^{V}),$ (4)
@@
where $\bm{Q}=\bm{F}_{\text{in}}$, $\bm{K}=\bm{f}^{1}_{\text{in}}$, $\bm{V}=\bm{f}^{1}_{\text{in}}$ are the query, key, and value, respectively, and $\bm{W}^{Q}$, $\bm{W}^{K}$, $\bm{W}^{V}$ are learnable matrices. The key and value are from the first frame of the video. Therefore, $\bm{S}$ represents the similarity between the query and the key from the first frame, which implicitly warps the first frame into subsequent frames [57]. Our technique simply applies the $\bm{S}$ computed from the original video when processing the augmented video, whose output features are used for noise prediction.